A body is falling free in a vacuum. The fall is necessarily related to the gravitational 

acceleration g and the height h from which the body is dropped. Use dimensional 

analysis to show that the velocity V of the falling body satisfies the relation 

V / √gh = constant


1
Expert's answer
2021-02-26T00:52:46-0500

Dimension of V is "\frac{m}{sec}."

Dimension of g is "\frac{m}{sec^2}."

Dimension of h is "m."

Thus, dimension of "\sqrt{gh}" is "\frac{m}{sec}=" dimension of V.

Therefore, the relation "\frac{V}{\sqrt{gh}}=constant."